OpenStudy (anonymous):

whats the y inter of y −x/3+ 2 = 0.

OpenStudy (karatechopper):

get y all by itself first. then u will know ur y intercept!

OpenStudy (anonymous):

Let's get y by itself y −x/3+ 2 = 0 add x/3 to both sides y + 2 = x/3 subtract 2 from both sides y = x/3 - 2 The number without the x term is the y-intercept. The answer is ...

OpenStudy (anonymous):

1/3 ?

OpenStudy (anonymous):

almost, the 1/3 is with the x or x-term this is called the slope . The answer is ...

OpenStudy (anonymous):

-2 <--------- answer

OpenStudy (anonymous):

ok thanks